# Universal Omni-Language Models

### v3

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. However, the capacity of these models to concurrently process and reason about multiple modalities remains underexplored, partly due to the lack of comprehensive modality-wise benchmarks. We introduce OmniBench, a novel benchmark designed to rigorously evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define language models capable of such tri-modal processing as the omni-language models (OLMs). OmniBench is distinguished by high-quality human annotations, ensuring that accurate responses require integrated understanding and reasoning across all three modalities. Our main findings reveal that: i) open-source OLMs exhibit critical limitations in instruction-following and reasoning capabilities within tri-modal contexts; and ii) most baselines models perform poorly (below 50% accuracy) even when provided with alternative textual representations of images or/and audio. These results suggest that the ability to construct a consistent context from text, image, and audio is often overlooked in existing MLLM training paradigms. To address this gap, we curate an instruction tuning dataset of 84.5K training samples, OmniInstruct, for training OLMs to adapt to multimodal contexts. We advocate for future research to focus on developing more robust tri-modal integration techniques and training strategies to enhance OLM performance across diverse scenarios. Codes, data and live leaderboard could be found at the project page.
@@ 1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly involving images, audio, and text (Li & Lu, 2024). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs that can handle two of the modalities, the ability to concurrently process and reason about the three aforementioned modalities remains a frontier yet to be fully explored.
@@
  • Compared to the open-source OLMs, the proprietary models perform better overall but suffer from more accuracy drops when ablating the image or audio input.
@@
These results underscore the importance of OmniBench as a tool for identifying areas of improvement and guiding research in multimodal systems. In the following sections, we 1) detail the data collection protocol of OmniBench; 2) present our evaluation results on current state-of-the-art MLLMs; 3) introduce the OmniInstruct dataset for omni-language model supervised fine-tuning; and 4) discuss the implications of our findings for the future of research and development. Through OmniBench, we aim to catalyze advancements in MLLMs, pushing the boundaries of artificial intelligence towards true omni-understanding capabilities.
2 Related Work
Multimodal Large Language Models.
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. In the audio domain, models like Whisper (Radford et al., 2022), BEATs (Chen et al., 2022), MERT (Li et al., 2023b), and CLAP (Wu et al., 2023b) have been developed as specialized encoders for speech, general audio, acoustic music, and music-text, respectively. These have been incorporated into more comprehensive systems such as SALMONN (Tang et al., 2023), LTU (Gong et al., 2023), Mu-llama, MusiLingo, and Audio-Flamingo. Notable progress in audio perception and instruction-following includes SALMONN (Tang et al., 2023), BLSPN (Wang et al., 2023a), Speech-LLaMAN (Wu et al., 2023a), and Qwen-Audio (Chu et al., 2023b), all demonstrating promising capabilities in audio-focused dialogues. In the visual domain, large visual language models have made significant strides, often leveraging pre-trained image encoders (Dosovitskiy, 2020; Touvron et al., 2020; Liu et al., 2021; Radford et al., 2021; Zhai et al., 2023). Notable examples include BLIP2 (Li et al., 2023a), which uses a Q-Former for visual-textual alignment, LLaVA (Liu et al., 2024b), which employs GPT-4 generated instruction data, and its successor LLaVA-Next (Liu et al., 2024a). Building on the LLaVA framework, models like QwenVL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and Yi-VL (Young et al., 2024) have achieved significant success through extensive pre-training. Despite these advancements, most existing MLLMs focus on a single modality for input processing while generating textual responses. While some models can process textual, audio, and visual inputs simultaneously, open-source models in this field generally exhibit less competitive capabilities compared to their closed-source counterparts. In this context, we define omni-language models (OLMs) as those capable of processing at least three different modalities of data simultaneously [^1] We target the models able to concurrently process image, audio, and text as a starting point since these are the most well-explored modalities in the field, but the ”omni” concept is extendable..
Multimodal Understanding Benchmark.
The vision-language benchmarks aim to test models’ ability to combine visual and language data in tasks like OCR, spatial awareness, multimodal information retrieval (e.g. SciMMIR (Wu et al., 2024a)), and multimodal reasoning skills. MM-Vet (Yu et al., 2023) focuses on visual question answering (VQA), requiring models to interpret visual data and respond to queries. MMBench (Liu et al., 2023c) evaluates models via multiple-choice tasks in both Chinese and English, covering diverse domains. MMStar (Chen et al., 2024a) conducts multi-task evaluations to test multimodal fusion capabilities. MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024) assess model performance on complex vision-language tasks, emphasizing sophisticated multimodal reasoning. MMRA (Wu et al., 2024b) is designed to evaluate the models’ multi-image relational association capability. In addition, there are several audio-understanding benchmarks. Aishell1 (Bu et al., 2017), Aishell2 (Du et al., 2018), and Librispeech (Panayotov et al., 2015) are designed for automatic speech recognition, while ClothoAQA targets audio QA tasks. For automatic audio captioning and vocal sound classification, researchers have curated Clotho (Drossos et al., 2020) and VocalSound (Gong et al., 2022). However, there is a significant lack of comprehensive understanding benchmarks to assess MLLMs’ ability to simultaneously process complementary information from the textual, audio, and visual inputs.
Audio-Visual Understanding Datasets.
In previous works on audio-visual question answering (AVQA), the focus has predominantly been on identifying visual objects, sounds, and their interrelations to foster multimodal understanding. For instance, the Pano-AVQA dataset (Yun et al., 2021) explores 360-degree panoramic video understanding with 5.4k videos and 51.7k QA pairs. However, it limits its scope to identifying existing objects or locations, omitting questions on causal inference and abstract concepts. Similarly, the Music-AVQA dataset (Li et al., 2022) comprises 9.3k videos, including 1.9k synthesized entries, and 45.9k QA pairs. These videos typically feature simple scenarios of music performances by one or two players or within a music chamber. The questions focus on existing objects, time and location, counting, and relationships but fail to address special symbols like music score images or causal inference. The Music-AVQA-2.0 dataset (Liu et al., 2024c) enhances Music-AVQA by collecting 1,230 manually curated musical ensemble performance videos and 8.1k newly created QA pairs designed to complement and diversify the original dataset, addressing biases in annotation and instrument types. However, it maintains the original types of QA pairs, not expanding into new categories of questions. The AVQA dataset (Yang et al., 2022), contains 57k QA pairs that do not necessarily require integration of both modalities for answering, illustrating its limitation in truly multimodal inquiry. For example, the presence of an auditory signal like a train whistle isn’t essential to deduce an action such as the lowering of a gear lever at a railroad crossing, suggesting that answers could be surmised from visual cues alone. This dataset includes questions on time, location, existing objects, causality, purpose, and counting, yet lacks coverage of actions, symbol concepts, and associations. VALOR dataset (Chen et al., 2023) is an audiovisual-language dataset designed for tri-modality model pre-training, comprising 1.18 M videos sourced and curated from AudioSet (Gemmeke et al., 2017). Recognizing captions derived from ASR or alt-texts fail to adequately align audio-language modalities; VALOR annotates audio-visual content by humans based on AudioSet tags to establish a clear correspondence among 3 modalities. VALOR is good for pre-training but does not include instruction following QA pairs. These gaps in current datasets underscore the need for more comprehensive AVQA datasets and evaluation benchmarks that can challenge and accurately measure a model’s capacity for deep multimodal integration and abstract reasoning, which are critical for advancing multimodal understanding in AI.
3 OmniBench
The OmniBench aims to create the first comprehensive benchmark for evaluating multimodal large language models that support simultaneous image, audio, and text inputs. While OmniBench is designed to evaluate the understanding capability of MLLMs on cross-modality complementary information, the models are required to interpret the multimodal input and provide accurate text answer. The problem could be formulated as following: given a tuple of (image, audio, text), the model is required to recognize the objects, re-build the contexts, and conduct reasoning based on the given information. The design logic and statistics of the dataset and the annotation protocols are introduced in this section.
Figure 2: The Taxonomy and Proportions of the 1142 Test Samples in OmniBench. The inner circle refers to the three major categories, and the traffic circle refers to the fine-grained task types.
3.1 Benchmark Design
Building on the foundation of existing multimodal benchmarks, our OmniBench introduces a refined taxonomy for task categorization that effectively captures a wide range of cognitive and reasoning abilities. As demonstrate in Figure 2, our framework organizes tasks into three primary categories: (1) (temporal)-spatial entity, which includes Object Identification for recognizing distinct entities and Contextual & Environmental for discerning the setting or backdrop of the events; (2) causal inference, comprised of Story Cause Description to infer narrative drivers, Current Action & Activity to understand ongoing dynamics, and Future Plot and Purpose Inference to anticipate subsequent developments; and (3) abstract concept, involving Identity & Relationship to identify and relate entities, Text & Symbols for symbolic interpretation, and Count & Quantity for numerical reasoning. This taxonomy is designed to evaluate both foundational perceptual skills and complex cognitive processes, thereby providing a comprehensive assessment of multimodal language models’ (MLLMs) abilities to integrate and interpret diverse information sources. OmniBench includes 1142 question-answer pairs, with details on task types, text length, and the characteristics of images and audio presented in Table 1. The audio content of the dataset is categorized into speech, sound events, and music, enriching the diversity of stimuli for evaluating the models’ tri-modal capabilities and aiding in the development of future omni-language models.
@@ 3.3 OmniInstruct
To construct a dataset that aligns with the challenges proposed in OmniBench, we enhanced each question to connect with an audio track and an image extracted from the corresponding video and filter it with VLMs for better quality. Notably, we avoid the first and last five frames of each video to exclude transitional or obscure incomplete scenes that might distort the task’s focus. For MSRVTT-QA train and valid subset, we discard videos without audio tracks and retain a dataset comprised 6,176 videos from the original set that include audio tracks alongside 151.7k QA pairs directly related to these videos. Then we use InternVL-2-76B to filter the questions from the three datasets mentioned above to delete (1) questions that can be answered only with an image, (2) questions irrelevant with image, potentially answerable only with audio, and (3) ambiguous or non-logical questions, where the detailed prompt and statistics could be found at Figure 6 and Table 7, Appendix B. Only 93k data samples remain for training and validation.
This curated dataset is essential for evaluating the nuanced capabilities of multimodal large language models to interpret and integrate multiple types of information, which is a first step towards enhancing reasoning performance and applicability in omni-modality scenarios.
@@ 4 Experiment Settings
We select three groups of MLLM baselines according to the modalities available: (i) omni-language models: MIO-Instruct (Wang et al., 2024b), AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b); (ii) vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (Wang et al., 2024a), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); (iii) audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023b), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not conclude them as in the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024..
@@
The main focus of OmniBench is to evaluate how well could the MLLMs understand and reconstruct the context given information from image ($I$), audio ($A$) and text ($T$) modalities. Setting up questions with four available options for the models, we use accuracy, i.e., the ratio matched letter of the correct option and model response, as the evaluation metric (n.b., the accuracy of a random guess model is $25\%$ under this setting). Additionally, we test the models in an ablation setting of removing one of the image or audio inputs to further reveal a more comprehensive reasoning capability of the baselines and verify the robustness of our benchmark. For baseline systems, please refer to section 4.
@@
5 Findings
5.1 Results on Omni-Language Models
Table 2 demonstrates that open-source omni-language model (OLM) baselines surpass random guessing accuracy across various settings. Notably, the UnifiedIO2 series demonstrates inconsistent performance scaling with model size, indicating challenges in effectively leveraging increased capacity for multimodal understanding.
Despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Besides, Video-Salmonn, developed by Bytedance, and Gemini, developed by Google, provide better results on music subsets compared to their performance on speech and music., potentially due to their large corpus of music videos, though the music ethics of training foundation models are still underdiscussion (Ma et al., 2024). Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-removed and audio-removed) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to open-source alternatives, its accuracy remains below 50%.
These findings underscore the challenging nature of OmniBench and highlight substantial room for improvement in multi-modal reasoning tasks. We anticipate the development of more competitive models on our benchmark in the near future, which will further advance the field of multi-modal AI.
@@ 5.1 Results on Omni-Language Models
Table 3: OLM Baselines Overall Results Grouped by Audio Type and Task Category. The accuracy numbers calculated by different audio types are at the upper table and the accuracy accross tasks categories are placed at the bottom table.

### v4

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have focused on integrating multiple modalities, yet their ability to simultaneously process and reason across different inputs remains underexplored. We introduce OmniBench, a novel benchmark designed to evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define language models capable of such tri-modal processing as omni-language models (OLMs). OmniBench features high-quality human annotations that require integrated understanding across all modalities. Our evaluation reveals that: i) open-source OLMs show significant limitations in instruction-following and reasoning in tri-modal contexts; and ii) most baseline models perform poorly (around 50% accuracy) even with textual alternatives to image/audio inputs. To address these limitations, we develop OmniInstruct, an 96K-sample instruction tuning dataset for training OLMs. We advocate for developing more robust tri-modal integration techniques and training strategies to enhance OLM performance. Codes and data could be found at our repo.
Machine Learning, ICML
@@ 1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly involving images, audio, and text (Li & Lu, 2024). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs handling two of the modalities, the ability to concurrently process and reason about the three modalities remains a frontier yet to be fully explored.
@@
  • In contrast, the proprietary models perform better overall but suffer from more accuracy drops when ablating the image or audio input.
  • Compared to models, the results of human evaluator show not only better overall performances but a distinct distribution (e.g., much better on ”Sound Event” audios and ”Abstract Concept” tasks).
@@
In the following sections, we 1) detail the data collection protocol of OmniBench; 2) present our evaluation results on current state-of-the-art MLLMs; 3) introduce the OmniInstruct dataset for omni-language model supervised fine-tuning; and 4) discuss the implications of our findings for the future of research and development.
@@
2 Related Work
Multimodal Large Language Models.
Recent advances in multimodal large language models have produced specialized encoders for audio processing (Radford et al., 2022; Chen et al., 2022; Li et al., 2023b; Wu et al., 2023b), which have been integrated into more comprehensive systems (Tang et al., 2023; Gong et al., 2023). Notable progress in audio-focused dialogues has demonstrated promising capabilities in instruction-following and perception (Tang et al., 2023; Wang et al., 2023a; Wu et al., 2023a; Chu et al., 2023b). In the visual domain, significant strides have been made through models that leverage pre-trained image encoders (Dosovitskiy, 2020; Touvron et al., 2020; Liu et al., 2021; Radford et al., 2021; Zhai et al., 2023). These models have achieved success through visual-textual alignment, GPT-4 generated instruction data, and extensive pre-training (Li et al., 2023a; Liu et al., 2024b, a; Bai et al., 2023; Wang et al., 2023b; Young et al., 2024). While most existing MLLMs focus on single-modality input processing, open-source models capable of processing multiple modalities generally show less competitive capabilities compared to closed-source alternatives. In this context, we define omni-language models (OLMs) as those capable of processing at least three different modalities simultaneously [^1] We target the models able to concurrently process image, audio, and text as a starting point since these are the most well-explored modalities in the field, but the ”omni” concept is extendable..
Multimodal Understanding Benchmark.
Current vision-language benchmarks evaluate various capabilities including OCR, spatial awareness, multimodal information retrieval, and reasoning skills (Wu et al., 2024a; Yu et al., 2023; Liu et al., 2023c; Chen et al., 2024a; Yue et al., 2024; Zhang et al., 2024; Wu et al., 2024b). These benchmarks assess models through multiple-choice tasks, complex vision-language problems, and multi-image relational understanding. In the audio domain, several benchmarks focus on speech recognition, audio QA tasks, and sound classification (Bu et al., 2017; Du et al., 2018; Panayotov et al., 2015; Drossos et al., 2020; Gong et al., 2022). However, there remains a significant gap in comprehensive benchmarks that can assess models’ ability to simultaneously process and integrate information from textual, audio, and visual inputs.
Audio-Visual Understanding Datasets.
Existing audio-visual question answering datasets have primarily focused on object and sound identification (Yun et al., 2021; Li et al., 2022; Liu et al., 2024c; Yang et al., 2022). While these datasets cover various aspects like temporal understanding and counting, they often lack comprehensive evaluation of causal inference and abstract reasoning. Some datasets don’t require true multimodal integration for answering questions, as responses can often be deduced from a single modality. Recent work has attempted to address these limitations through human annotation and improved alignment between modalities (Chen et al., 2023; Gemmeke et al., 2017), but still lacks instruction-following evaluation capability. These limitations highlight the need for more comprehensive datasets that can effectively assess models’ multimodal integration and reasoning abilities.
@@ 2 Related Work
3 OmniBench
The OmniBench aims to create a comprehensive benchmark for evaluating multimodal large language models that support simultaneous image, audio, and text inputs. While OmniBench is designed to evaluate the understanding capability of MLLMs on cross-modality complementary information, the models are required to interpret the multimodal input and provide accurate text answer. The problem could be formulated as following: given a tuple of (image, audio, text), the model is required to recognize the objects, re-build the contexts, and conduct reasoning based on the given information. The design logic and statistics of the dataset and the annotation protocols are introduced in this section.
3.1 Benchmark Design
Building on the foundation of existing multimodal benchmarks, our OmniBench introduces a refined taxonomy for task categorization that effectively captures a wide range of cognitive and reasoning abilities. Our framework organizes tasks into three primary categories: (1) (temporal)-spatial entity, which includes Object Identification for recognizing distinct entities and Contextual & Environmental for discerning the setting or backdrop of the events; (2) causal inference, comprised of Story Cause Description to infer narrative drivers, Current Action & Activity to understand ongoing dynamics, and Future Plot and Purpose Inference to anticipate subsequent developments; and (3) abstract concept, involving Identity & Relationship to identify and relate entities, Text & Symbols for symbolic interpretation, and Count & Quantity for numerical reasoning. This taxonomy is designed to evaluate both foundational perceptual skills and complex cognitive processes, thereby providing a comprehensive assessment of multimodal language models’ (MLLMs) abilities to integrate and interpret diverse information sources. OmniBench includes 1142 question-answer pairs, with details on task types, text length, and the characteristics of images and audio presented in Table 1. The audio content of the dataset is categorized into speech, sound events, and music, enriching the diversity of stimuli for evaluating the models’ tri-modal capabilities and aiding in the development of future omni-language models.
@@ 3.3 OmniInstruct
Table 2: Summarization of OmniInstruct Data Distribution across Modalities and Question Types.
image audio how many what who when&where others
Train 84,580 - 1,131 56,168 22,449 959 3,873
Valid 11,525 - 80 4,498 1,884 66 4,997
Total 96,105 - 1,211 60,666 24,333 1,025 8,870
To construct a dataset that aligns with the challenges proposed in OmniBench, we enhanced each question to connect with an audio track and an image extracted from the corresponding video and filter it with VLMs for better quality. Notably, we avoid the first and last five frames of each video to exclude transitional or obscure incomplete scenes that might distort the task’s focus. For MSRVTT-QA train and valid subset, we discard videos without audio tracks and retain a dataset comprised 6,176 videos from the original set that include audio tracks alongside 151.7k QA pairs directly related to these videos. Then we use InternVL-2-76B to filter the questions from the three aforementioned datasets to delete (1) questions that can be answered only with an image, (2) questions irrelevant with image, potentially answerable only with audio, and (3) ambiguous or non-logical questions, where the detailed prompt and statistics could be found at Figure 5 and Table 9, Appendix B. After the processing pipeline, only 96k data samples remain for training and validation.
@@ 4 Experiment Settings
We select three groups of MLLM baselines according to the modalities available: (i) omni-language models: MIO-Instruct (Wang et al., 2024b), AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b), the VITA series (Fu et al., 2024, 2025), OpenOmni (Luo et al., 2025), Baichuan-Omni-1.5 (Li et al., 2025), Qwen-2.5-Omni (Xu et al., 2025); (ii) vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (Wang et al., 2024a), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); (iii) audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023b), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not conclude them as in the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024.. Besides, we invite three musician with higher education background to test on the benchmark and use the average accuracy as a human expert baseline.
@@
The main focus of OmniBench is to evaluate how well could the MLLMs understand and reconstruct the context given information from image ($I$), audio ($A$) and text ($T$) modalities. Setting up questions with four available options for the models, we use accuracy, i.e., the ratio matched letter of the correct option and model response, as the evaluation metric (n.b., the accuracy of a random guess model is $25\%$ under this setting). Additionally, we test the models in an ablation setting of removing one of the image or audio inputs to further reveal a more comprehensive reasoning capability of the baselines and verify the robustness of our benchmark.
@@
Human Evaluator 63.19% - -
Table 4: OLM Baselines Overall Results Grouped by Audio Type.
@@
Human Evaluator 58.71% 75.85% 64.15%
Table 5: OLM Baselines Overall Results Grouped by Task Category.
@@
Human Evaluator 72.91% 55.51% 53.87% 71.41% 65.48% 59.38% 45.33% 66.67%
5 Findings
We evaluate the selected baseline systems in multiple meticulously designed settings, and provide insights on the development of OLMs based on the results. All the claims stated in this section are supported by statistic significance verification, for more information please refer to Appendix C.
5.1 Results on Omni-Language Models
Overall
Table 3 demonstrates that most open-source OLM baselines surpass random guessing accuracy across various setting, and the latest developed one (Qwen-2.5-Omni) could even outperform the proprietary models. Notably, the UnifiedIO2 series demonstrates inconsistent performance scaling with model size, indicating challenges in effectively leveraging increased capacity for multimodal understanding. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-removed and audio-removed) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to most of the open-source alternatives, its accuracy remains below 50%.